3. How many possible of size n = 3 can be 

drawn from a population of size 12?

a. 36 c. 144

b. 1728 d. 220

4.Taken from the same population, which 

sample size will give a smaller standard error 

of the mean?

a. 35 c. 18

b. 12 d. 9

6. What is the finite population correction

factor if the size of the population is 200 and 

the sample size is 25?

a. 0.979 c. 0.938

b. 0.879 d. 0.856

8. If a population has a mean of 23.7, what is 

the mean of the sampling distribution?

a. less than 23.7

b. larger than 23.7

c. closer to 23.7

d. the same as 23.7

9. If 50 randomly selected college students 

take the examination, what is the probability 

that the mean time it takes the group to 

complete the test will be less than 43 

minutes?

a. 23% c. 4.24%

b. 0.23% d. 0.42%

10. Please refer to item 9.

 Does it seem reasonable that a college 

student would finish the examination in less 

than 43 minutes?

a. YES b. NO


1
Expert's answer
2021-06-17T14:46:56-0400

3. "\dbinom{12}{3}=\dfrac{12!}{3!(12-3)!}=220"

d. 220


4. "SE=\sigma/\sqrt{n}"

a. 35


6.


"\dfrac{N-n}{N-1}=\dfrac{200-25}{200-1}\approx0.879"

b. 0.879 


8.

d. the same as 23.7


9.


"P(Z<\dfrac{43-46.2}{8/\sqrt{50}})\approx0.0023"

b. 0.23%


10.

b. NO